# Solving Dicke superradiance analytically: A compendium of methods

### v1

@@ IV A probabilistic approach
The constraint $j_{N}+\ldots+j_{m}+(N-m)=j$ ensures that the final time is $t=j\Delta t$ and $(N-m)$ jumps have occurred resulting in the state |$\ket{m}$. The numbers $j_{N},\ldots,j_{m}$ count the number of time steps the system stays in the corresponding state between jumps. Notice that the product over jumps leads to $h_{N}\cdots h_{m+1}\Gamma^{N-m}(\Delta t)^{N-m}$ and we can write
@@
$\rho_{m}(t)=\sum_{p\in\text{poles}}\text{Res}\Big{[}z^{j}\Big{(}\frac{h_{N}\Gamma\Delta t}{z-s_{N}}\Big{)}\cdots\Big{(}\frac{h_{N}\Gamma\Delta t}{z-s_{m}}\Big{)}\Big{]}_{z=p}.$ (32)
Notice that when evaluating $z$ at the poles, the differences in the denominator are $s_{k}-s_{k^{\prime}}=-\Gamma\Delta t(h_{k}-h_{k^{\prime}})$ canceling the terms containing $\Gamma\Delta t$ in the numerator and also leading to a term $(-1)^{N-m}$. In the limit of infinitesimal time steps, $\Delta t\to 0$, indicates that $j=t/\Delta t$ can be taken to infinity. In this limit, any term $s_{k}^{j}=(1-h_{k}\Gamma t/j)^{j}$ can be approximated by its limit value $e^{-\Gamma h_{k}t}$. The expression above then again maps fully into the solution presented in Sec. I.
@@ V Rate equations approach
$\rho_{m}(t)=\sum_{j=m}^{N}(-1)^{N-m}h_{N}\cdots h_{m+1}\prod_{\begin{subarray}{c}j^{\prime}=m\\ j^{\prime}\neq j\end{subarray}}^{N}\frac{1}{h_{j}-h_{j^{\prime}}}e^{-h_{j}\Gamma t},$ (35)
@@ VI.1 Time evolution operator
where we see that the top right element is not a pure exponential decay. These decays of the form $\Gamma te^{-h_{m}\Gamma t}$ stem from the occurring EP2222s, ultimately resulting in superradiance.

### v3

@@ IV A probabilistic approach
The constraint $j_{N}+\ldots+j_{m}+(N-m)=j$ ensures that the final time is $t=j\Delta t$ and that $(N-m)$ jumps have occurred in this time interval, leading to the target state |$\ket{m}$. The numbers $j_{N},\ldots,j_{m}$ count the number of time steps the system stays in the corresponding state between jumps. Notice that the product over jumps leads to $h_{N}\cdots h_{m+1}\Gamma^{N-m}(\Delta t)^{N-m}$ and we can write
@@
$\rho_{m}(t)=\sum_{p\in\text{poles}}\text{Res}\Big{[}z^{j}\Big{(}\frac{h_{N}\Gamma\Delta t}{z-s_{N}}\Big{)}\cdots\Big{(}\frac{h_{m}\Gamma\Delta t}{z-s_{m}}\Big{)}\Big{]}_{z=p}.$ (34)
Notice that when evaluating $z$ at the poles, the differences in the denominator are $s_{k}-s_{k^{\prime}}=-\Gamma\Delta t(h_{k}-h_{k^{\prime}})$ canceling the terms containing $\Gamma\Delta t$ in the numerator and also leading to a term $(-1)^{N-m}$. In the limit of infinitesimal time steps, $\Delta t\to 0$, $j=t/\Delta t$ can be taken to infinity. In this limit, any term $s_{k}^{j}=(1-h_{k}\Gamma t/j)^{j}$ can be approximated by its limit value $e^{-\Gamma h_{k}t}$. The expression above then again maps fully into the solution presented in Sec. I.
@@ V Rate equations approach
$\rho_{m}(t)=\sum_{j=m}^{N}(-1)^{N-m}h_{N}\cdots h_{m+1}\Bigg{[}\prod_{\begin{subarray}{c}j^{\prime}=m\\ j^{\prime}\neq j\end{subarray}}^{N}\frac{1}{h_{j}-h_{j^{\prime}}}\Bigg{]}e^{-h_{j}\Gamma t},$ (37)
@@ VI.1 Time evolution operator
where we see that the top right element is not a pure exponential decay. These decays of the form $\Gamma te^{-h_{m}\Gamma t}$ stem from the occurring EP2222s.